# RectifID: Personalizing Rectified Flow with Anchored Classifier Guidance

### v1

@@ 1 Introduction
Existing personalization methods are limited in these two aspects, as they require an extra finetuning or pre-training stage. For example, the pioneering works (Gal et al., 2023; Ruiz et al., 2023a) finetune conditional embeddings or model parameters per subject, resulting in suboptimal efficiency and identity consistency due to lack of domain knowledge. On the other hand, the recently prevailing tuning-free methods (Wei et al., 2023; Ye et al., 2023; Li et al., 2024; Wang et al., 2024) pre-train a conditioning adapter to encode subject features into the generation process. However, their models must be pre-trained on extensive domain-specific data, e.g. LAION-Face 50M (Zheng et al., 2022), which is costly in the first place, and cannot be transferred flexibly across different data domains, e.g. from human faces to common live subjects and objects, and even to multiple subjects.
@@ 2 Background
Personalized image generation studies incorporating user-specified subjects into the text-to-image generation pipeline. To preserve the subject’s identity, the seminal works Textual Inversion (Gal et al., 2023) and DreamBooth (Ruiz et al., 2023a) finetune conditional embeddings or model parameters for each subject, which imposes high computational costs. Subsequent literature resorts to more efficient parameters (Hu et al., 2022; Han et al., 2023; Yuan et al., 2023) or a pre-trained subject encoder (Wei et al., 2023; Ye et al., 2023) to allow personalization within a few minutes or even without tuning. At the other end, a recent trend is the reuse of existing discriminators to improve identity consistency, such as extracting discriminative face features as the condition (Ye et al., 2023; Wang et al., 2024) or as a training objective for the encoder (Peng et al., 2024; Gal et al., 2024; Guo et al., 2024). However, these models require extensive pre-training on domain-specific data, e.g. LAION-Face 50M (Zheng et al., 2022). In contrast, our method is a training-free approach that exploits existing discriminators based on the recent rectified flow model, allowing flexible personalization for a variety of tasks.
@@
Recently, rectified flow has shown promising efficiency (Liu et al., 2024) and quality (Esser et al., 2024; Yan et al., 2024) in text-to-image generation. Our work extends its capabilities and theoretical properties to personalized image generation via classifier guidance.
@@ 4.1 Experimental Settings
Metrics. Three metrics are considered: identity similarity, prompt consistency, and computation time. The first two are measured using an ArcFace model (Deng et al., 2019) and CLIP encoders (Radford et al., 2021), while the latter is tested on an NVIDIA A800 GPU. We reproduce the latest methods IP-Adapter (Ye et al., 2023), PhotoMaker (Li et al., 2024), and InstantID (Wang et al., 2024) for a comprehensive comparison, and also include the existing baselines in Pang et al. (2024).
@@
Table 1: Quantitative comparison for face-centric personalization. The computation time is measured in seconds on an NVIDIA A800. Unlike the previous state-of-the-art methods that require training on large face datasets (the number of images is listed for reference), our method achieves superior performance in a training-free manner, by exploiting the guidance from an off-the-shelf discriminator.
@@
InstantID (Wang et al., 2024) SDXL 60M 0.5806 0.3071 6
@@
Figure 3: Qualitative comparison for face-centric personalization. See Figs. 7 and 8 for more samples.
@@ 4.2 Main Results
For face-centric qualitative comparison in Fig. 3, our method remains advantageous as its generated images by the guidance of the face discriminator exhibit high identity consistency. In comparison, InstantID (Wang et al., 2024) delivers a near level of consistency by controlling face landmarks, but sometimes distorts the face shape (the first and third images) and contains much less natural variation. More generated samples are provided in Figs. 7 and 8 in the appendix.
@@ 4.3 Ablation Study
To justify the effectiveness of our proposed classifier guidance, Figs. 6 and 2 compare it with two variants: the previously derived guidance without anchor, namely using Eq. 8, and a gradient descent method on the initial noise similar to DOODL (Wallace et al., 2023) and D-Flow (Ben-Hamu et al., 2024). The figure depicts that the gradient descent is unstable (left) and converges relatively slowly (right) despite using momentum and $\ell_{2}$ regularization. And its identity preservation is sensitive to the learning rate. Though our new fixed-point formulation allows for a more stable layout, the initial version fails to converge as the face feature keeps drifting. In contrast, our full method exhibits better stability (left) and faster convergence (right) by implicitly regularizing the flow trajectory to be close and straight. This is further supported by the quantitative comparison, where our method delivers better identity and prompt consistency than the alternatives. Additional analysis for hyperparameter sensitivity is provided in Section D.2.
@@ Appendix B Related Work on Classifier Guidance
Since the proposal of classifier guidance [Dhariwal and Nichol, 2021] which uses a special noise-aware classifier as training-free guidance for diffusion models, many efforts have been made in order to extend its applicability to off-the-shelf loss guidance. They can be grouped into three categories: (1) Early literature focuses on simpler objectives for linear inverse problems such as image super-resolution, deblurring, and inpainting [Chung et al., 2022, 2023, Wang et al., 2023, Zhu et al., 2023]. (2) These methods can be extended to more complex discriminators through various approximations. Yu et al. [2022] and Song et al. [2023] use Tweedie’s formula and Monte Carlo method, respectively, to estimate the integrated classifier guidance. Bansal et al. [2023] and He et al. [2024] perform updates directly in the clean data space, with the latter imposing an additional manifold constraint. (3) A recent line of work directly uses gradient descent with specific diffusion models. To enable their gradient computation, DiffusionCLIP [Kim et al., 2022a] relies on shortened ODE trajectories, while FlowGrad [Liu et al., 2023b] adopts a non-uniform ODE discretization and decomposed gradient computation. DOODL [Wallace et al., 2023], DNO [Karunratanakul et al., 2024] and D-Flow [Ben-Hamu et al., 2024] use invertible models or flow models to backpropagate gradient to the initial noise.Our proposed method is related to the third category, focusing on rectified flow whose approximation remains understudied. Moreover, it features a fixed-point formulation with a convergence guarantee for ideal rectified flow, which allows potentially better stability over the existing approaches, e.g. gradient descent on initial noise, as empirically validated in Section 4.3.
@@ Appendix C Experimental Settings
For face-centric personalization, our method is implemented with the antelopev2 model pack from the InsightFace library. [^1] https://github.com/deepinsight/insightface Specifically, it detects and crops the face regions with an SCRFD model [Guo et al., 2022], and then extracts face features using an ArcFace model [Deng et al., 2019] trained on Glint360K [An et al., 2021], consistent with most personalization methods that use a face model [Ye et al., 2023, Wang et al., 2024, Gal et al., 2024]. The resulting face feature is compared with the reference image to compute a cosine loss, which serves as the classifier guidance signal.
@@
Evaluation. For face-centric personalization, we follow Pang et al. [2024] in evaluating on the first 200 images in the CelebA-HQ dataset [Liu et al., 2015, Karras et al., 2018] with 20 text prompts including 15 realistic prompts and 5 stylistic prompts. The evaluation process reuses the code from Celeb Basis [Yuan et al., 2023], [^2] This is mentioned at https://github.com/lyuPang/CrossInitialization#metrics. which first detects the face region using a PIPNet [Jin et al., 2021] with a threshold of 0.5 and then computes the cosine similarity on face features extracted by an ArcFace model [Deng et al., 2019]. It should be noted that our method adopts a different face detector and different alignment and cropping methods, so it does not overfit the evaluation protocol. For the baselines, in addition to those compared in Pang et al. [2024], we also evaluate the recently proposed IP-Adapter [Ye et al., 2023], PhotoMaker [Li et al., 2024], and InstantID [Wang et al., 2024] on their recommended settings. The number of their sampling steps is set to 30 for a fair comparison. The checkpoint version of IP-Adapter is ip-adapter-full-face_sd15. The image size is set to 512$\times$512 for IP-Adapter and 1024$\times$1024 for PhotoMaker and InstantID based on SDXL [Podell et al., 2024]. In the qualitative analysis, we also include Celeb Basis [Yuan et al., 2023] as a baseline method.
@@
Licenses. The piecewise rectified flow [Yan et al., 2024] used in the main experiments is released under the BSD-3-Clause License and the 2-rectified flow [Liu et al., 2023a, 2024] used in Section D.3 is released under the MIT License. The InsightFace library for face detection and recognition is released under the MIT License, while its pre-trained models are available for non-commercial research purposes only. The OWL-ViT [Minderer et al., 2022] and DINOv2 [Oquab et al., 2023] models for object detection and feature extraction are released under the Apache-2.0 License. For evaluation, the code of Celeb Basis [Yuan et al., 2023] is licensed under the MIT License. The CelebA-HQ [Liu et al., 2015, Karras et al., 2018] and DreamBooth [Ruiz et al., 2023a] datasets for quantitative and qualitative evaluation are released under the CC BY-NC 4.0 and CC-BY-4.0 licenses.
@@ Appendix D Additional Results
D.1 More Visualizations
More examples of our generated image are provided in Figs. 7, 8, 9 and 10. For face-centric personalized image generation, it is shown that our method can follow a variety of text prompts to generate both realistic or stylistic images while preserving the user-specified identity from a diverse group of people. Although there exist minor differences in the person’s age and hairstyles, the face looks very similar to the reference image. For subject-driven generation, new results are presented from additional subject types, including different breeds of cats and dogs, and some regularly shaped objects such as vases, demonstrating the flexibility of our approach across different use cases. For multi-subject generation, it naturally blends multiple human faces or live subjects into a single image while maintaining the visual quality and semantics, revealing a wider range of potentially interesting applications. Overall, our method demonstrates to be effective and flexible for various personalization tasks.
D.2 Ablation Study
@@
D.3 Experiments with 2-Rectified Flow
Figures 12 and 13 present additional qualitative results on a vanilla 2-rectified flow [Liu et al., 2023a, 2024] finetuned on Stable Diffusion 1.4 [Rombach et al., 2022]. As can be seen, our method continues to deliver satisfactory identity preservation when moving to a different model, despite a noticeable drop in the generation quality. Concretely, it integrates target subjects with some quite challenging prompts, such as a person swimming or getting a haircut, while showing very little interference with the original background, e.g. jungles and cityscapes, validating the effectiveness of our method.
Note that we also experimented with $K=1$ on a single-step InstaFlow [Liu et al., 2024] distilled from this 2-rectified flow, but found that it tended to converge to slightly distorted images. This may be attributed to the larger modeling error inherent in InstaFlow’s distillation process, which reduces the effectiveness of our approach assuming each flow trajectory segment is well-trained and straight.
@@ Appendix E Broader Impacts and Limitations
Broader impacts. The proposed method can be integrated with the emerging rectified flow-based models to enhance identity preservation and versatile control over existing AI art production pipelines. However, as a training-free personalization technique, it may increase the risk of image faking and have negative societal effects. Some immediate remedies include text-based prompt filters and AI-generated image detection, but it remains an open problem for a more principled solution, for which we advocate further research on data watermarking and model unlearning as potential mitigations.
@@
Figure 7: Additional face-centric personalization results with piecewise rectified flow [Yan et al., 2024], which is based on Stable Diffusion 1.5 [Rombach et al., 2022]. Our method achieves high identity consistency. See Fig. 12 for results on the vanilla 2-rectified flow [Liu et al., 2023a, 2024].
@@
Figure 8: Additional face-centric personalization results with piecewise rectified flow [Yan et al., 2024], which is based on Stable Diffusion 1.5 [Rombach et al., 2022]. Our method achieves high identity consistency. See Fig. 12 for results on the vanilla 2-rectified flow [Liu et al., 2023a, 2024].
@@
Figure 9: Additional subject-driven generation results with piecewise rectified flow [Yan et al., 2024], which is based on Stable Diffusion 1.5 [Rombach et al., 2022]. Our approach preserves the identity of both live subjects and some regularly shaped objects. Please see Fig. 13 for more examples using the vanilla 2-rectified flow [Liu et al., 2023a, 2024].
@@
Figure 12: Face-centric personalization results with vanilla 2-rectified flow [Liu et al., 2023a, 2024], which is based on Stable Diffusion 1.4 [Rombach et al., 2022]. Our method preserves their identities well while remaining faithful to the text prompt during generation.
@@
Figure 13: Subject-driven generation results with vanilla 2-rectified flow [Liu et al., 2023a, 2024], which is based on Stable Diffusion 1.4 [Rombach et al., 2022]. Examples for additional categories of cats, dogs, and objects are included to demonstrate the effectiveness of our approach.

### v2

@@ 1 Introduction
Existing personalization methods are limited in these two aspects, as they require an extra finetuning or pre-training stage. For example, the pioneering works (Gal et al., 2023; Ruiz et al., 2023a) finetune conditional embeddings or model parameters per subject, resulting in suboptimal efficiency and identity consistency due to lack of domain knowledge. On the other hand, the recently prevailing tuning-free methods (Wei et al., 2023; Ye et al., 2023; Li et al., 2024; Wang et al., 2024b) pre-train a conditioning adapter to encode subject features into the generation process. However, their models must be pre-trained on extensive domain-specific data, e.g. LAION-Face 50M (Zheng et al., 2022), which is costly in the first place, and cannot be transferred flexibly across different data domains, e.g. from human faces to common live subjects and objects, and even to multiple subjects.
@@ 2 Background
Personalized image generation studies incorporating user-specified subjects into the text-to-image generation pipeline. To preserve the subject’s identity, the seminal works Textual Inversion (Gal et al., 2023) and DreamBooth (Ruiz et al., 2023a) finetune conditional embeddings or model parameters for each subject, which imposes high computational costs. Subsequent literature resorts to more efficient parameters (Hu et al., 2022; Han et al., 2023; Yuan et al., 2023) or a pre-trained subject encoder (Wei et al., 2023; Ye et al., 2023) to allow personalization within a few minutes or even without tuning. At the other end, a recent trend is the reuse of existing discriminators to improve identity consistency, such as extracting discriminative face features as the condition (Ye et al., 2023; Wang et al., 2024b) or as a training objective for the encoder (Peng et al., 2024; Gal et al., 2024; Guo et al., 2024). However, these models require extensive pre-training on domain-specific data, e.g. LAION-Face 50M (Zheng et al., 2022). In contrast, our method is a training-free approach that exploits existing discriminators based on the recent rectified flow model, allowing flexible personalization for a variety of tasks.
@@
Recently, rectified flow has shown promising efficiency (Liu et al., 2024b) and quality (Esser et al., 2024; Yan et al., 2024) in text-to-image generation. Our work extends its capabilities and theoretical properties to personalized image generation via classifier guidance.
@@ 4.1 Experimental Settings
Metrics. Three metrics are considered: identity similarity, prompt consistency, and computation time. The first two are measured using an ArcFace model (Deng et al., 2019) and CLIP encoders (Radford et al., 2021), while the latter is tested on an NVIDIA A800 GPU. We reproduce the latest methods IP-Adapter (Ye et al., 2023), PhotoMaker (Li et al., 2024), and InstantID (Wang et al., 2024b) for a comprehensive comparison, and also include the existing baselines in Pang et al. (2024).
@@
Table 1: Quantitative comparison for face-centric personalization. The inference time is measured in seconds on an NVIDIA A800. Unlike the previous state-of-the-art methods that require training on large face datasets (the number of images is listed for reference), our method achieves superior performance in a training-free manner, by exploiting the guidance from an off-the-shelf discriminator.
@@
InstantID (Wang et al., 2024b) SDXL 60M 0.5806 0.3071 6
@@
Figure 3: Qualitative comparison for face-centric personalization. See Figs. 11, 12, 9 and 10 for more samples.
@@ 4.2 Main Results
For face-centric qualitative comparison in Fig. 3, our method remains advantageous as its generated images by the guidance of the face discriminator exhibit high identity consistency. In comparison, InstantID (Wang et al., 2024b) delivers a near level of consistency by controlling face landmarks, but sometimes distorts the face shape (the first and third images) and contains much less natural variation. More generated samples are provided in Figs. 11 and 12 in the appendix.
@@ 4.3 Ablation Study
To justify the effectiveness of our proposed classifier guidance, Figs. 6 and 2 compare it with two variants: the previously derived guidance without anchor, namely using Eq. 8, and a gradient descent method on the initial noise similar to DOODL (Wallace et al., 2023) and D-Flow (Ben-Hamu et al., 2024). The figure depicts that the gradient descent is unstable (left) and converges relatively slowly (right) despite using momentum and $\ell_{2}$ regularization. And its identity preservation is sensitive to the learning rate. Though our new fixed-point formulation allows for a more stable layout, the initial version fails to converge as the face feature keeps drifting. In contrast, our full method exhibits better stability (left) and faster convergence (right) by implicitly regularizing the flow trajectory to be close and straight. This is further supported by the quantitative comparison, where our method delivers better identity and prompt consistency than the alternatives. Further analysis for hyperparameter sensitivity is provided in Section D.3.
4.4 Generalization
(a) Segmentation map
(b) Style transfer
Figure 7: Experimental results for more controllable generation tasks. The first column shows the guidance, and the rest are the generated results. Our method is extended to various controllable generation tasks by incorporating the guidance functions from Universal Guidance (Bansal et al., 2024).
To validate the generalizability of our approach to broader application scenarios, we have extended it to more controllable generation tasks by directly using the guidance functions from Universal Guidance (Bansal et al., 2024). The experimental results under the guidance of segmentation map or style image are illustrated in Fig. 7. As shown, our classifier guidance can perform both tasks without any additional tuning, faithfully following the various forms of control signals provided by the user. This confirms the adaptability of our approach for various controllable generation tasks. Additional generalization analysis of our method for broader diffusion models is presented in Section D.1.
@@ Appendix B Related Work on Classifier Guidance
Since the proposal of classifier guidance [Dhariwal and Nichol, 2021] which uses a special noise-aware classifier as training-free guidance for diffusion models, many efforts have been made in order to extend its applicability to off-the-shelf loss guidance. They can be grouped into three categories: (1) Early literature focuses on simpler objectives for linear inverse problems such as image super-resolution, deblurring, and inpainting [Chung et al., 2022, 2023, Wang et al., 2023, Zhu et al., 2023]. (2) These methods can be extended to more complex discriminators through various approximations. Yu et al. [2022], Song et al. [2023] use Tweedie’s formula and Monte Carlo method, respectively, to estimate the integrated classifier guidance. Bansal et al. [2024], He et al. [2024] perform updates directly in the clean data space, with the latter imposing an additional manifold constraint. Similarly, Gaussian spherical constraint is explored in Yang et al. [2024]. Mou et al. [2024] advance these techniques to more versatile editing tasks. The above methods are further unified in Ye et al. [2024]. (3) A recent line of work directly uses gradient descent with specific diffusion models. To enable their gradient computation, DiffusionCLIP [Kim et al., 2022a] relies on shortened ODE trajectories, while FlowGrad [Liu et al., 2023b] adopts a non-uniform ODE discretization and decomposed gradient computation. DOODL [Wallace et al., 2023], DNO [Karunratanakul et al., 2024] and D-Flow [Ben-Hamu et al., 2024] use invertible models or flow models to backpropagate gradient to the initial noise.Our proposed method is related to the third category, focusing on rectified flow whose approximation remains understudied. Moreover, it features a fixed-point formulation with a convergence guarantee for ideal rectified flow, which allows potentially better stability over the existing approaches, e.g. gradient descent on initial noise, as empirically validated in Section 4.3.
@@ Appendix C Experimental Settings
For face-centric personalization, our method is implemented with the antelopev2 model pack from the InsightFace library. [^1] https://github.com/deepinsight/insightface Specifically, it detects and crops the face regions with an SCRFD model [Guo et al., 2022], and then extracts face features using an ArcFace model [Deng et al., 2019] trained on Glint360K [An et al., 2021], consistent with most personalization methods that use a face model [Ye et al., 2023, Wang et al., 2024b, Gal et al., 2024]. The resulting face feature is compared with the reference image to compute a cosine loss, which serves as the classifier guidance signal.
@@
Evaluation. For face-centric personalization, we follow Pang et al. [2024] in evaluating on the first 200 images in the CelebA-HQ dataset [Liu et al., 2015, Karras et al., 2018] with 20 text prompts including 15 realistic prompts and 5 stylistic prompts. The evaluation process reuses the code from Celeb Basis [Yuan et al., 2023], [^2] This is mentioned at https://github.com/lyuPang/CrossInitialization#metrics. which first detects the face region using a PIPNet [Jin et al., 2021] with a threshold of 0.5 and then computes the cosine similarity on face features extracted by an ArcFace model [Deng et al., 2019]. It should be noted that our method adopts a different face detector and different alignment and cropping methods, so it does not overfit the evaluation protocol. For the baselines, in addition to those compared in Pang et al. [2024], we also evaluate the recently proposed IP-Adapter [Ye et al., 2023], PhotoMaker [Li et al., 2024], and InstantID [Wang et al., 2024b] on their recommended settings. The number of their sampling steps is set to 30 for a fair comparison. The checkpoint version of IP-Adapter is ip-adapter-full-face_sd15. The image size is set to 512$\times$512 for IP-Adapter and 1024$\times$1024 for PhotoMaker and InstantID based on SDXL [Podell et al., 2024]. In the qualitative analysis, we also include Celeb Basis [Yuan et al., 2023] as a baseline method.
@@
Licenses. The piecewise rectified flow [Yan et al., 2024] used in the main experiments is released under the BSD-3-Clause License and the 2-rectified flow [Liu et al., 2023a, 2024b] used in Section D.4 is released under the MIT License. The InsightFace library for face detection and recognition is released under the MIT License, while its pre-trained models are available for non-commercial research purposes only. The OWL-ViT [Minderer et al., 2022] and DINOv2 [Oquab et al., 2023] models for object detection and feature extraction are released under the Apache-2.0 License. For evaluation, the code of Celeb Basis [Yuan et al., 2023] is licensed under the MIT License. The CelebA-HQ [Liu et al., 2015, Karras et al., 2018] and DreamBooth [Ruiz et al., 2023a] datasets for quantitative and qualitative evaluation are released under the CC BY-NC 4.0 and CC-BY-4.0 licenses.
@@ Appendix D Additional Results
D.1 Generalization
Input painting in Van Gogh style reading on the train concert poster Greek sculpture faded film in the snow
(a) SD-Turbo
Input stained glass window pointillism painting oil painting sketch in a police outfit manga drawing
(b) Phased consistency model
Figure 8: Generalization to few-step diffusion models, including SD-Turbo [Sauer et al., 2024] and phased consistency model [Wang et al., 2024a], both distilled from SD and using 4 sampling steps in inference. The results show that our method is effective for personalizing broader diffusion models.
While our classifier guidance is derived based on rectified flow, the same idea can be generalized to some few-step diffusion models by assuming straightness of their trajectories within each time step. We empirically demonstrate this in Fig. 8 with two popular few-step diffusion models, SD-Turbo [Sauer et al., 2024] and phased consistency model [Wang et al., 2024a]. As the results indicate, our method effectively personalizes these diffusion models to generate identity-preserving images. We will continue to explore this approach for other generative models in future research.
D.2 More Visualizations
More examples of our generated image are provided in Figs. 11, 12, 9, 10, 13 and 14. For face-centric personalized image generation, it is shown that our method can follow a variety of text prompts to generate both realistic or stylistic images while preserving the user-specified identity from a diverse group of people. Although there exist minor differences in the person’s age and hairstyles, the face looks very similar to the reference image. In particular, the method demonstrates good generalizability among different piecewise rectified flows based on SDXL [Podell et al., 2024] and SD 1.5 [Rombach et al., 2022]. For subject-driven generation, new results are presented from additional subject types, including different breeds of cats and dogs, and some regularly shaped objects such as vases, demonstrating the flexibility of our approach across different use cases. For multi-subject generation, it naturally blends multiple human faces or live subjects into a single image while maintaining the visual quality and semantics, revealing a wider range of potentially interesting applications. Overall, our method demonstrates to be effective and flexible for various personalization tasks.
D.3 Ablation Study
@@
D.4 Experiments with 2-Rectified Flow
Figures 16 and 17 present additional qualitative results on a vanilla 2-rectified flow [Liu et al., 2023a, 2024b] finetuned on Stable Diffusion 1.4 [Rombach et al., 2022]. As can be seen, our method continues to deliver satisfactory identity preservation when moving to a different model, despite a noticeable drop in the generation quality. Concretely, it integrates target subjects with some quite challenging prompts, such as a person swimming or getting a haircut, while showing very little interference with the original background, e.g. jungles and cityscapes. These results clearly validate the effectiveness of our proposed method in alternative rectified flow models.
Note that we also experimented with $K=1$ on a single-step InstaFlow [Liu et al., 2024b] distilled from this 2-rectified flow, but found that it tended to converge to slightly distorted images. This may be attributed to the larger modeling error inherent in InstaFlow’s distillation process, which reduces the effectiveness of our approach assuming each flow trajectory segment is well-trained and straight.
@@ Appendix E Broader Impacts and Limitations
Broader impacts. The proposed method can be integrated with the emerging rectified flow-based models to enhance identity preservation and versatile control over existing AI art production pipelines. However, as a training-free personalization technique, it may increase the risk of image faking and have negative societal effects. Some immediate remedies include text-based prompt filters and AI-generated image detection, but it remains an open problem for a more principled solution, for which we advocate further research on data watermarking and model unlearning as potential mitigations. To further clarify it, we provide a more detailed explanation of these mitigations below:
  • Prompt filtering, model unlearning: Since our method keeps the diffusion model intact, existing techniques for regulating diffusion models can be applied seamlessly, including prompt filters or unlearning methods. The former can be applied explicitly like the text filters in SD models, or implicitly via CFG as in Schramowski et al. [2023]. The latter approaches involve finetuning the diffusion model to remove the ability to generate harmful content [Gandikota et al., 2023, Kumari et al., 2023].
  • Data watermarking: To prevent misuse of personal images, one could add a protective watermark to their images [Van Le et al., 2023, Liu et al., 2024a]. With this watermark or perturbation, the image can no longer be learned by common personalization methods. However, it is unclear how robust these watermarks are to training-free methods such as ours. An alternative watermarking scheme is to embed special watermark to the images generated by our proposed model, which would be invisible to the users yet identifiable by us (i.e., the tech provider). Images with such watermarks will be marked as being artificial.
  • AI-generated image detection: As a post-hoc safety measure, it helps to distinguish fake images generated by the attackers from real images. Beyond above watermark-based scheme, more sophisticated data-driven methods have attracted increasing interest from the AI community. Despite that current methods still lack accuracy, we believe that developing reliable and widely available AI-generated image detectors is an important research direction.
@@
Another important issue is the lack of pre-trained discriminators. To address this in the short term, we suggest first training a specialized discriminator and then applying our classifier guidance. There are two reasons for doing this instead of finetuning the generator directly: (1) training/finetuning a discriminator is usually more efficient and stable than training/finetuning a generator; (2) it can take full advantage of domain images that have no captions or even labels by using standard contrastive learning loss. In the future, scaling up vision-language models may be a general solution for these domains. The current models such as GPT-4V [OpenAI, 2023] have demonstrated certain generalizability across visual understanding tasks. As they continue to improve in generalizability and robustness, they will become a viable source for guiding diffusion models in new domains.
Input graduating after finishing PhD wear a magician hat and a blue coat in a garden as White Queen wearing headphones concert poster pencil drawing
Input in a police outfit holding roses in front of the Eiffel Tower wear a magician hat and a blue coat in a garden reading on the train buckled in his seat on a plane sipping coffee at a cafe
Input stained glass window painting in Van Gogh style watercolor painting colorful mural on a street wall driving a car buckled in his seat on a plane
Input cubism painting wearing a brown jacket and a hat, holding a whip ink painting graduating after finishing PhD cooking a gourmet meal in blue suit pencil drawing
Input typing a paper on a laptop piloting a fighter jet wearing a hat holding roses in front of the Eiffel Tower concert poster in a chef outfit
Input cave mural piloting a fighter jet wear a magician hat and a blue coat in a garden driving a car swimming pencil drawing
Input as an amazon warrior wear a magician hat and a blue coat in a garden in a cowboy hat wearing a hat watercolor painting pencil drawing
Figure 9: Additional face-centric personalization results with piecewise rectified flow [Yan et al., 2024] based on SDXL [Podell et al., 2024]. Our method is compatible with more advanced base models and provides sophisticated personalization results.
Input Banksy art piloting a fighter jet in a cowboy hat colorful mural on a street wall concert poster watercolor painting
Input pointillism painting piloting a fighter jet wearing a hat faded film in a police outfit playing the LEGO toys
Input cubism painting piloting a fighter jet buckled in his seat on a plane wear a magician hat and a blue coat in a garden in a police outfit watercolor painting
Input surrounded by bookshelves wearing a scifi spacesuit in space reading on the train coding in front of a computer swimming in a chef outfit
Input pointillism painting in a cowboy hat wearing a hat faded film concert poster sipping coffee at a cafe
Input stained glass window piloting a fighter jet western vibes, sunset, rugged landscape holding roses in front of the Eiffel Tower concert poster wearing a sunglass on a boat
Input colorful mural on a street wall wearing a scifi spacesuit in space western vibes, sunset, rugged landscape wear a magician hat and a blue coat in a garden as White Queen reading on the train
Figure 10: Additional face-centric personalization results with piecewise rectified flow [Yan et al., 2024] based on SDXL [Podell et al., 2024]. Our method is compatible with more advanced base models and provides sophisticated personalization results.
@@
Figure 11: Additional face-centric personalization results with piecewise rectified flow [Yan et al., 2024], which is based on Stable Diffusion 1.5 [Rombach et al., 2022]. Our method achieves high identity consistency. See Fig. 16 for results on the vanilla 2-rectified flow [Liu et al., 2023a, 2024b].
@@
Figure 12: Additional face-centric personalization results with piecewise rectified flow [Yan et al., 2024], which is based on Stable Diffusion 1.5 [Rombach et al., 2022]. Our method achieves high identity consistency. See Fig. 16 for results on the vanilla 2-rectified flow [Liu et al., 2023a, 2024b].
@@
Figure 13: Additional subject-driven generation results with piecewise rectified flow [Yan et al., 2024], which is based on Stable Diffusion 1.5 [Rombach et al., 2022]. Our approach preserves the identity of both live subjects and some regularly shaped objects. Please see Fig. 17 for more examples using the vanilla 2-rectified flow [Liu et al., 2023a, 2024b].
@@
Figure 16: Face-centric personalization results with vanilla 2-rectified flow [Liu et al., 2023a, 2024b], which is based on Stable Diffusion 1.4 [Rombach et al., 2022]. Our method preserves their identities well while remaining faithful to the text prompt during generation.
@@
Figure 17: Subject-driven generation results with vanilla 2-rectified flow [Liu et al., 2023a, 2024b], which is based on Stable Diffusion 1.4 [Rombach et al., 2022]. Examples for additional categories of cats, dogs, and objects are included to demonstrate the effectiveness of our approach.